OpenStudy (anonymous):

can anyone help me to integrate this guestion INTEGRATION OF (x^2)/(1+x^6) help me to solve this...

hartnn (hartnn):

u= x^3 du=... ?

hartnn (hartnn):

also, you'll need, \(\large \int \dfrac{1}{1+u^2}du=\tan^{-1}u+c\)

OpenStudy (anonymous):

ok i get it now..thanks..

hartnn (hartnn):

welcome ^_^